# chattersPwCHparticipants with chatbot history

### v1

@@ Abstract.
Programming students have a widespread access to powerful Generative AI tools like ChatGPT. While this can help understand the learning material and assist with exercises, educators are voicing more and more concerns about an over-reliance on generated outputs and lack of critical thinking skills. It is thus important to understand how students actually use generative AI and what impact this could have on their learning behavior. To this end, we conducted a study including an exploratory experiment with 37 programming students, giving them monitored access to ChatGPT while solving a code understanding and improving exercise. While only 23 of the students actually opted to use the chatbot, the majority of those eventually prompted it to simply generate a full solution. We observed two prevalent usage strategies: to seek knowledge about general concepts and to directly generate solutions. Instead of using the bot to comprehend the code and their own mistakes, students often got trapped in a vicious cycle of submitting wrong generated code and then asking the bot for a fix. Those who self-reported using generative AI regularly were more likely to prompt the bot to generate a solution. Our findings indicate that concerns about potential decrease in programmers’ agency and productivity with Generative AI are justified. We discuss how researchers and educators can respond to the potential risk of students uncritically over-relying on generative AI. We also discuss potential modifications to our study design for a large-scale replications.
@@ 1. Introduction
With the public release of tools such as OpenAI Codex (Chen et al., 2021), ChatGPT (OpenAI, 2022) and GitHub Copilot (GitHub, 2022), students in programming courses now have access to code authoring tools capable of solving coding tasks and exam questions entirely (Berrezueta-Guzman and Krusche, 2023; Chen et al., 2021; Finnie-Ansley et al., 2022, 2023; Savelka et al., 2023). However, such tools also tend to confidently present incorrect information (Cao et al., 2021; Kabir et al., 2024; Tian et al., 2024) or generate subtly incorrect code (Dakhel et al., 2023) which may be difficult to detect for beginners (Popovici, 2023; Stanik et al., 2018). Due to the relative novelty of these tools, ChatGPT for example having been released in November 2022, their impact is not yet fully understood, with ongoing research into applications and challenges. Nonetheless, the tools are already being used by numerous students (Prather et al., 2023a). This increasingly requires educators to respond.
A recent interview study with individual students revealed interest in using the technology for purposes like generating supplementary learning materials, but also skipping coursework they don’t find engaging enough (Zastudil et al., 2023). Meanwhile, university instructors are largely unsure about how many of their students are using ChatGPT and to what extent (Prather et al., 2023a). Across multiple surveys and position papers, educators have expressed concerns, particularly about over-reliance on generated content and about the ease of cheating (Debby R. E. Cotton and Shipway, 2023; Kasneci et al., 2023; Lau and Guo, 2023; Prather et al., 2023a; Zastudil et al., 2023).
In response, even educators with a generally positive sentiment towards have begun implementing restrictions and bans in the classroom. Long term, two major approaches appear to emerge from this development: either a) embracing as a tool not only for teaching but also as a core software development tools, or b) preventing and restricting its use (Lau and Guo, 2023). Both approaches involve considerable work, as the curriculum and course design need to be modified to either include or exclude these tools from the classroom and the students’ overall learning journey.
Before taking such impactful decisions, it is particularly important to understand how programming students are actually using tools and whether the usage strategies would primarily help them learn or rather avoid coursework. This work takes the first step into this direction. We report on a study with programming beginners who had optional monitored access to ChatGPT. We observed and analyzed their interactions while solving programming task, focusing on the following research questions:
  • RQ1: Could a student pass an introductory programming course using only ChatGPT-generated answers?
@@
  • RQ3: How much work within a programming exercise do students delegate to ChatGPT (i.e. how much autonomous thinking do they keep)?
The remainder is structured as follows: Section 2 describes the study design including research methods, participants, and tasks. Section 3 presents the results along the RQs. Then, Section 4 summarizes the findings, highlights a few observations, and discusses the study implications and limitations. Finally, Section 5 discusses related work and Section 6 concludes the paper.
@@ 2.1. Research Methodology
To answer RQ1, we evaluated the task-solving performance of GPT-3.5 and GPT-4 on all exercise assignments of a large first-semester introductory programming course. To answer RQ2 and RQ3, we conducted an experimental study with students in the same course, who volunteered to solve an exercise while having access to a monitored ChatGPT interface. We recorded and labeled the students’ interactions with ChatGPT, and performed pattern analysis on the resulting chat logs.
@@ 2.1.1. Task-Solving Evaluation
For all assignments that must to be solved to pass the programming course, we created individual prompts to be submitted to the GPT API. Some exercises are split into subtasks with different levels of granularity. We separated these tasks into minimal congruent subtask groups, i.e. subtasks that shared the same project context, such that only immediately related subtasks were joined into a single prompt. Additionally, if the exercise required knowledge of an existing template project for students to build off of, that project was also appended to the prompt text. The full input to the language model consisted of a short system prompt instructing it to solve a Java programming exercise, the exercise text as given to students, and – if applicable – a code snippet containing the project context. We evaluated GPT-3.5 using the gpt-3.5-turbo-1106 model, and GPT-4 using the gpt-4-1106-preview model. The temperature parameter for both models was set to 0.0 to reduce randomness in the output. We inserted the generated solution into the appropriate project files, performed the usual tests tutors do for the exercises to pass, and recorded the results alongside the generated code.
As the course relies heavily on manual reviews by course tutors, any submitted code not only needs to meet the functional requirements of the exercises, but also match the paradigms and coding conventions taught in the course. In fact, the majority of in-person exercises included in the studied course did not have automated unit tests to verify the student solutions. Instead, in addition to code reviews, course tutors performed a series of manual tests, by asking students to demonstrate specific behaviors using the interactive BlueJ object interface. These required code reviews and manual tests were specified in the tutor instructions for each exercise.
@@
  • Course Requirements. The answer must fulfill the requirements shared across all exercises of the course, most importantly the adherence to the course’s coding conventions (Style). These are not explicitly stated in each question text, but still expected of the students.
  • Authenticity. A student must have plausibly been able to derive (and explain) the provided answer from the course’s learning material (Learn). Students utilizing more advanced concepts must demonstrate a strong understanding of their solution, otherwise their answer is considered likely plagiarized and thus rejected. Additionally, any comments and explanations must be consistent with the code they describe (Intent), or the authenticity of the presented solution is called into question.
@@ 2.1.2. Student Experiment
In the experimental study, we presented students with an optional code authoring exercise that they were asked to solve to the best of their ability. Additionally, participants were given access to a ChatGPT-like chatbot to assist them with the exercise. The use of the bot was not required. We recorded participants code changes, chatbot conversations, copy/paste events, and their solution submissions. The exercise was realized as a quiz activity on the Moodle platform, with two introductory self-assessment Likert scale questions and a CodeRunner coding exercise using the same configuration as described in the next section. Students were used to this e-learning platform as about 20% of their usual assignments were handled there.
The chatbot UI featured a subset of the functionality of ChatGPT. Participants could create conversation threads and hold a back-and-forth conversation with the chatbot. Responses were gradually revealed as they were generated, and rendered with full Markdown support. Overall the bot very closely mimics the look of ChatGPT as a generic tool. Figure 1 shows a screenshot of the chatbot. It uses the GPT API to generate responses to participants’ prompts. In the experiment we used the gpt-3.5-turbo-1106 model, which was the latest publicly available at the time.
@@
Our data analysis includes a quantitative and a qualitative component. We collected aggregate activity metrics on the students’ attempts, performance, interactions, and the timeline of events. Additionally, two authors manually labeled independently each student prompt and each chatbot response to differentiate between types of requested information and identify interaction patterns.
With RQ2, we set out to understand of the kinds of prompts students would submit to an AI assistant. For this, we coded the general objective of each prompt based on labels listed in Table 1. The labels were initially created through a deductive approach (Neuendorf, 2017), and then consolidated them after reviewing the actual experiment chat logs. Initially, we had outlined a label Retry for participants requesting a corrected solution specifically without describing the preceding failure, i.e. asking the to simply “try again”. We found this to be a very rare occurrence and very similar to the label Fix. Therefore we decided to merge both labels. For coarser analysis, we also grouped multiple prompt labels into a category Codegen, which represents all prompts asking for code output from the . Similarly, the category Support represents prompts asking for knowledge or assistance in understanding or completing the exercise. These categories are also shown in Table 1.
Table 1. Labels for participant-submitted prompts, with descriptions and examples.
@@
Solve Request a solution to the problem directly, including by simply relying the question text.
@@
Example: “[question code] Which of these methods might produce a null-related error?”
@@
We also coded the chatbot responses following the labels listed in Table 2. These labels are non-exclusive, i.e. a response can simultaneously be assigned two labels. For example, a generated code solution that is additionally explained in text would be assigned both Solves and Explains.
@@
Solves Provides a code snippet that attempts to solve one of the exercise problems [^1] We additionally determined programmatically whether a chatbot response contained any code, including generic example snippets..
@@
Lastly, for both generated factual claims and code solutions, we labeled whether they were correct or incorrect; for solution attempts, this meant whether the code fulfilled the stated requirements. In cases where it was ambiguous whether the response was correct given the intent of the request, such as when the preceding prompt did not contain any instructions, neither label was applied.
To answer RQ3, inspired by the concerns of “blindly copying and pasting solutions” (Prather et al., 2023a, p.15), we tracked clipboard events – i.e. cut, copy and paste actions – both on the chatbot and exercise pages. To trace the flow of information, we then grouped the recorded actions into pairs of a cut or copy event directly followed by a paste event, and compared the text contents to ensure the events in the pair were actually related.
@@
Additionally, we analyzed the similarity between the current and previous submission, as well as to any code generations between two submissions. However, this measure does not differentiate between functional and non-functional modifications. For example, changing a local variable name has the same impact on similarity as changing out a method call, despite the former not having any impact on the solution. To address this issue, we manually reviewed all code diffs generated by consecutive submissions with a code generation in between, and labeled them according to the labels listed in Table 3.
@@
We used the NOSEP algorithm (Wu et al., 2017b) for this analysis, as we were looking for consecutive, non-overlapping subsequences that may occur multiple times in one sequence. We specified a gap range of $[0,2]$ to allow for minor tolerance against actions such as Others prompts within a pattern, and a minimum support of $10$. To identify each pattern occurrence and link it back to the participant, we then used the NETLAP algorithm (Wu et al., 2017a) on each of the reported frequent patterns.
@@ 2.2. Setting and Participants
Both the task-solving evaluation and the experiment were conducted as part of the introductory course Softwareentwicklung 1 (SE1) at the University of Hamburg. The course typically accommodates between 500 and 700 students each year. Acceptance into the course is not contingent on prior software development experience and, as such, the course is designed to accommodate programming novices. The course takes place over a 14-week semester, covering the basics of programming, control flow and logic, object-oriented programming, data structures, and code quality.
Each week, a worksheet with 2-3 major exercises is released to students, which they solve through in-person pair programming during two hour long lab sessions. These exercises can be code authoring, answering questions, or constructing diagrams. The exercises are inspired by Barnes and Kölling (Barnes and Kölling, 2009).
Upon completing a major exercise, students request a course tutor to review their work and sign off on it. During this review, tutors will ask additional questions to test the students’ understanding of the material, allowing them to “identify and correct misunderstandings early and provide immediate feedback with personalized explanations” (Häring and Maalej, 2019). Additionally, tutors can verify the authenticity of the presented work in this process. Tutors are instructed to reject work that either does not meet the course requirements or that the students cannot adequately explain. While this does not prevent plagiarism outright, students are disincentivized from presenting plagiarized work unless they have a thorough understanding of the concepts and reasoning behind it.
An interactive quiz hosted on the Moodle platform is released alongside the worksheet each week. The course requires students to achieve a grade of at least 90% averaged across all quizzes. Students can attempt the quiz as often as they like, but only within a week of its release. The assessments also include code authoring questions using the Moodle CodeRunner (Lobb and Harlow, 2016) plugin. In these questions, students are given the problem statement and a minimal code editor in their browser. The student code is automatically compiled and tested whenever the student uses the ”Check” button to submit an attempt. Immediate feedback from the test runner is provided, showing compilation errors, runtime exceptions or test failures. There is no penalty for repeated attempts.
Experimental Task

### v2

# chattersCUchatbot-using participant

@@ Abstract.
Programming students have a widespread access to powerful Generative AI tools like ChatGPT. While this can help understand the learning material and assist with exercises, educators are voicing more and more concerns about an overreliance on generated outputs and lack of critical thinking skills. It is thus important to understand how students actually use generative AI and what impact this could have on their learning behavior. To this end, we conducted a study including an exploratory experiment with 37 programming students, giving them monitored access to ChatGPT while solving a code authoring exercise. The task was not directly solvable by ChatGPT and required code comprehension and reasoning. While only 23 of the students actually opted to use the chatbot, the majority of those eventually prompted it to simply generate a full solution. We observed two prevalent usage strategies: to seek knowledge about general concepts and to directly generate solutions. Instead of using the bot to comprehend the code and their own mistakes, students often got trapped in a vicious cycle of submitting wrong generated code and then asking the bot for a fix. Those who self-reported using generative AI regularly were more likely to prompt the bot to generate a solution. Our findings indicate that concerns about potential decrease in programmers’ agency and productivity with Generative AI are justified. We discuss how researchers and educators can respond to the potential risk of students uncritically over-relying on Generative AI. We also discuss potential modifications to our study design for large-scale replications.
@@
copyright: ccdoi: 10.1145/3715762journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE045publicationmonth: 7ccs: Social and professional topics Software engineering educationccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
With the public release of tools such as ChatGPT (OpenAI, 2022) and GitHub Copilot (GitHub, 2022), students in programming courses now have access to code authoring tools capable of solving coding exercises (Chen et al., 2021; Berrezueta-Guzman and Krusche, 2023; Savelka et al., 2023) and exam questions entirely (Finnie-Ansley et al., 2023). However, such tools also tend to confidently present incorrect information (Cao et al., 2021; Kabir et al., 2024; Tian et al., 2024) or generate subtly incorrect code (Dakhel et al., 2023) which may be difficult to detect for beginners (Popovici, 2023; Stanik et al., 2018). Furthermore, reliance on code generation can negatively impact code authoring skills (Kazemitabaar et al., 2023; Jošt et al., 2024).
Due to the relative novelty of these tools, ChatGPT for example released in November 2022, their effects are not yet fully understood. Initial studies have shown conflicting results regarding the impact of on learners, with some reporting neutral or slightly positive changes (Kazemitabaar et al., 2023; Vadaparty et al., 2024; Xue et al., 2024), while others even found adverse effects (Jošt et al., 2024; Prather et al., 2024). Nonetheless, the widespread use of these tools among students (Prather et al., 2023a) increasingly requires educators to respond.
A recent interview study with individual students revealed interest in using the technology for purposes like generating supplementary learning materials, but also skipping coursework they don’t find engaging enough (Zastudil et al., 2023). Meanwhile, university instructors are largely unsure about how many of their students are using ChatGPT and to what extent (Prather et al., 2023a). Across multiple surveys and position papers, educators have expressed concerns, particularly about over-reliance on generated content and about the ease of cheating (Cotton et al., 2024; Prather et al., 2023a; Zastudil et al., 2023; Becker et al., 2023).
In response, even educators with a positive sentiment towards have begun implementing restrictions and bans in the classroom. Two major approaches appear to emerge from this development: either a) embracing as a tool not only for teaching but also as a core software development tool, or b) preventing and restricting its use (Lau and Guo, 2023). Both approaches involve considerable work, as the curriculum and course design need to be modified to either include or exclude these tools from the classroom and the students’ overall learning journey.
Before taking such impactful decisions, it is particularly important to understand how programming students are actually using tools and whether the usage strategies would primarily help them learn or rather avoid coursework. To this end, we report on a study with programming beginners who had monitored access to ChatGPT, focusing on the following research questions:
  • RQ1: Could a student pass our introductory programming course using only ChatGPT-generated answers?
@@
  • RQ3: How much work do students delegate to ChatGPT and how much autonomous thinking do they keep within a programming exercise?
Our study is unique concerning the task we gave to participants, since it was not directly solvable by ChatGPT. Students had to comprehend and reason about the code (Maalej et al., 2014) to arrive at the solution or to formulate a relevant prompt for getting effective assistance from the bot. By analyzing the interactions of participants with the bot as well as their previous experience, we observed a widespread “lazy” cycle of asking the bot to solve the task, executing its wrong or misleading suggestion, and then copying the error message and asking again. By analyzing the diff of the submitted code, we found that students relying on the bot were rather following the bot suggestions even with completely different solution approaches, while students without the bot were rather trying to incrementally improve their solutions. In 62% of cases with generations, the modified submission was more similar to the generated code than to the previous submission. We report on a batch of qualitative and quantitative analyses and discuss potential implications on educational assessment as well as effective usage of by programming students and novice developers.
The remainder of the paper is structured as follows: Section 2 describes the design of our study to answer these RQs, including research methods, participants, and tasks. Section 3 presents the results along the RQs. Then, Section 4 summarizes the findings, highlights a few observations, and discusses the study implications and limitations. Finally, Section 5 discusses related work while Section 6 concludes the paper.
@@ 2. Study Design
We first present the research methodology for the task-solving evaluation and student experiment. Then, we introduce the study setting, including the programming course and the participants.
@@ 2.1. Research Methodology
To answer RQ1, we evaluated the task-solving performance of GPT models on all exercise assignments of our large first-semester introductory programming course. To answer RQ2 and RQ3, we conducted an experimental study with students in the same course, who volunteered to solve an exercise while having access to a monitored ChatGPT interface. We recorded and labeled the students’ interactions with ChatGPT, and performed pattern analysis on the resulting chat logs.
@@ 2.1.1. Task-Solving Evaluation
For all assignments that must to be solved to pass the programming course, we created individual prompts to be submitted to the GPT API. Some exercises are split into subtasks with different levels of granularity. We separated these tasks into minimal congruent subtask groups, i.e. subtasks that shared the same project context, such that only immediately related subtasks were joined into a single prompt. Additionally, if the exercise required knowledge of an existing template project for students to build off of, that project was also appended to the prompt text. The full input to the language model consisted of a short system prompt instructing it to solve a Java programming exercise, the exercise text as given to students, and – if applicable – a code snippet containing the project context.
At time of the study, GPT-4 was only available through paid access and with strict rate limitations within ChatGPT. We thus decided to evaluate both the latest model and the one most accessible to students. For GPT-3.5, this was the model gpt-3.5-turbo-1106. For GPT-4, we used the gpt-4-1106-preview model. The temperature parameter for both models was set to 0.0 to reduce randomness in the output. We inserted the generated solution into the appropriate project files, performed the usual tests human tutors do for the exercises to pass, and recorded the results alongside the generated code.
As the course relies heavily on manual reviews by course tutors (Häring and Maalej, 2019), any submitted code not only needs to meet the functional requirements of the exercises, but also match the paradigms and coding conventions taught in the course. In fact, the majority of in-person exercises included in the studied course did not have automated unit tests to verify the student solutions. Instead, in addition to code reviews, course tutors performed a series of manual tests, by asking students to demonstrate specific behaviors using the interactive BlueJ object interface. These code reviews and manual tests, which are required to pass, are specified in the tutor instructions for each exercise.
@@
  • Course Requirements. The answer must fulfill the requirements shared across all exercises of the course, most importantly the adherence to the course’s coding conventions (Style). These are not explicitly stated in each question text, but still expected from the students.
  • Authenticity. A student must plausibly be able to derive (and explain) the provided answer from the learning material (Learn). Students utilizing more advanced concepts must demonstrate a strong understanding of their solution, otherwise their answer is considered likely plagiarized and thus rejected. Additionally, any comments and explanations must be consistent with the code they describe (Intent), or the authenticity of the presented solution is questioned.
@@ 2.1.2. Student Experiment
In the experimental study, we presented students with an optional code understanding and authoring exercise that they were asked to solve to the best of their ability. Additionally, participants were given access to a ChatGPT-like chatbot to assist them with the exercise. The use of the bot was not required. We recorded participants code changes, chatbot conversations, copy/paste events, and their solution submissions. The exercise was realized as a quiz activity on the Moodle e-learning platform, with two introductory self-assessment semantic scale questions and a CodeRunner coding task. Students were used to this e-learning platform as about 20% of their usual assignments were handled there.
The chatbot UI featured a subset of the functionality of ChatGPT. Participants could create conversation threads and hold a back-and-forth conversation with the chatbot. Responses were gradually revealed as they were generated, and rendered with full Markdown support. Overall the bot very closely mimics the look of ChatGPT as a generic tool. Figure 1 shows a screenshot of the chatbot. It uses the GPT API to generate responses to participants’ prompts. In the experiment, we used the gpt-3.5-turbo-1106 model, which was the latest publicly available at the time.
@@
Our data analysis includes a quantitative and a qualitative component. We collected aggregate activity metrics on the students’ attempts, performance, interactions, and the timeline of events. Additionally, two authors manually labeled independently each student prompt and each chatbot response to identify types of requested information and interaction patterns.
With RQ2, we set out to understand the kinds of prompts students submit to an AI assistant. For this, we coded the general objective of each prompt based on labels listed in Table 1. The labels were initially created through a deductive approach (Maalej and Robillard, 2013; Neuendorf, 2017), and then consolidated after reviewing the actual chat logs. After independent labeling, the first annotator reviewed all disagreements and resolved those where one label was clearly inapplicable by definition. The remaining disagreements were individually discussed and mutually resolved by the annotators. For the prompt labels, the annotators had an initial agreement of 72% ($\kappa=0.65$).
Initially, we had outlined a label Retry for participants requesting a corrected solution specifically without describing the preceding failure, i.e. asking the bot to simply “try again”. We found this to be a very rare occurrence and very similar to the label Fix. Therefore, we decided to merge both labels. For coarser analysis, we also grouped multiple prompt labels into a category Codegen, which represents all prompts asking for code output. Similarly, the category Support represents prompts asking for knowledge or assistance in understanding or completing the exercise.
Table 1. Labels for prompts submitted by participants, with descriptions and examples.
@@
Solve Request a solution to the problem directly, including by pasting the question text.
@@
Example: “[question code] Which of these methods might produce a null error?”
@@
We also coded the chatbot responses following the labels listed in Table 2. These labels are non-exclusive, i.e. a response can simultaneously be assigned two labels. For example, a generated code solution that is additionally explained in text would be assigned both Solves and Explains. After independent labeling, the annotators had 80% agreement on the labels for this task.
@@
Solves Provides a code snippet that attempts to solve one of the exercise problems.
@@
Lastly, for both generated factual claims and code solutions, we labeled whether they were correct or incorrect. For solution attempts, this means whether the code fulfills the stated requirements. In cases where it was ambiguous whether the response was correct given the intent of the request, such as when the preceding prompt did not contain any instructions, neither label was applied.
To answer RQ3, inspired by the concerns of “blindly copying and pasting solutions” (Prather et al., 2023a, p.15), we tracked clipboard events – i.e. cut, copy and paste actions – both on the chatbot and exercise pages. To trace the flow of information, we then grouped the recorded actions into pairs of a cut or copy event directly followed by a paste event. We compared the text contents to ensure the events in the pair were actually related.
@@
Due to the small space of possible solutions, even semantically heavily modified submissions would have some similarity to previously generated code. Through manual review, we found that submissions with a similarity score above 90% were unambiguously close, structurally and semantically, to the generated code solution. We thus consider this our threshold for a close match.
Additionally, we analyzed the similarity between the current and previous submission, as well as to any code generations between two submissions. However, this measure does not differentiate between functional and non-functional modifications. For example, changing a local variable name has the same impact on similarity as changing out a method call, despite the former not having any impact on the solution. To address this issue, we manually reviewed all code diffs generated by consecutive submissions with a code generation in between, and labeled them according to the labels listed in Table 3. The agreement on this labeling task was 82% ($\kappa=0.72$).
@@
For this analysis, we used the NOSEP algorithm by Wu et al. (2017b), as we were looking for consecutive, non-overlapping subsequences that may occur multiple times in one sequence. We specified a gap range of $[0,2]$ to allow for minor tolerance against actions such as Others prompts within a pattern, and a minimum support of $10$. To identify each pattern occurrence and link it back to the participant, we used the NETLAP algorithm (Wu et al., 2017a) on each mined pattern.
@@ 2.2. Setting and Participants
2.2.1. Course Structure
Both the task-solving evaluation and the experiment were conducted as part of the introductory course Softwareentwicklung 1 (SE1) at the University of Hamburg. The course typically accommodates between 500 and 700 students each year. The course is designed to accommodate programming novices without any prior experience. The course takes place over a 14-week semester, covering the basics of programming, control flow and logic, object-oriented programming, data structures, and code quality.
Each week, a worksheet with 2-3 major exercises is released to students, which they solve through in-person pair programming during two-hour lab sessions. The exercises include code authoring, answering questions, or creating diagrams. The exercises are inspired by Barnes and Kölling (2009).
Upon completing a major exercise, students request a tutor to review and accept their work. During this review, tutors ask additional questions to test the students’ understanding of the material, allowing them to “identify and correct misunderstandings early and provide immediate feedback with personalized explanations” (Häring and Maalej, 2019). Additionally, tutors can verify the authenticity of the presented work. Tutors are instructed to reject work that either does not meet the course requirements or that the students cannot adequately explain. While this does not prevent plagiarism outright, students are disincentivized from presenting plagiarized work unless they have a thorough understanding of the concepts and reasoning behind it.
An interactive quiz hosted on the Moodle platform is released alongside the worksheet each week. The course requires students to achieve a grade of at least 90% averaged across all quizzes. Students can attempt the quiz as often as they like, but only within a week of its release. The assessments also include code authoring questions using the Moodle CodeRunner (Lobb and Harlow, 2016) plugin. In these questions, students get the problem statement and a minimal code editor in their browser. Their code is automatically compiled and tested whenever they use the ”Check” button to submit an attempt. Immediate feedback from the test runner is provided, showing compilation errors, runtime exceptions, or test failures. There is no penalty for repeated attempts.
2.2.2. Experimental Task